OpenStudy (anonymous):

Help please. Will give medal and fan. The function f(x) = 396(3)^x represents the growth of a dragonfly population every year in a remote swamp. Pierre wants to manipulate the formula to an equivalent form that calculates every month, not every year. Which function is correct for Pierre's purposes? A) f(x) = 33(3)^x B) f(x) = 396(3^12)^x/12 C) f(x) = 396(3^1/12)^12x D) f(x) = 4752(3)^x

OpenStudy (anonymous):

I was thinking the answer would be B but not sure

OpenStudy (linn99123):

I would honestly say C would be the correct answer to me... You would square it not divide it :)

OpenStudy (anonymous):

oh ok well i started to say B or C but then second guessed myself math is not my best subject. thank you

OpenStudy (anonymous):

B and C are the same aquations?

OpenStudy (anonymous):

396(3)^x = 396(3^12)^x/12 = 396(3^12)^x/12

OpenStudy (anonymous):

So are you saying that either B or C would be correct?

OpenStudy (linn99123):

B and C are different B) f(x) = 396(3^12)^x/12 you divide x by 12 C) f(x) = 396(3^1/12)^12x you multiply x while squared i believe.

OpenStudy (anonymous):

the answer is A

OpenStudy (anonymous):

396*3^x is the growth in one year. then (396*3^x)/12 is the growth in a month

OpenStudy (anonymous):

(396*3^x)/12=33*3^x

OpenStudy (anonymous):

oh ok thank you so much

OpenStudy (linn99123):

Dont forget to state weather its right or wrong :D It helps out future users!